Maria's brother Joe is three years less than twice Maria's age. Their combined ages add up to 24. How old is Maria?
Answer by addingup(3677)

J = 2M-3
J+M = 24 substitute for J:
2M-3+M = 24 regroup
2M+M-3 = 24
3M = 27
M = 9
and Joe = 2M-3 = 2(9)-3 = 18-3 = 15
----------------------------
check:
9+15 = 24 Correct